# Forcing Planets to Evolve: The Relationship Between Uranus and Neptune at Late Stages of Dynamical Evolution

### v1

@@ Abstract
The dynamical properties of the bodies in the outer Solar System hold information regarding the planets’ orbital histories. In early Solar System numerical simulations, where chaos is a primary driver, it is difficult to explore parameter space in a systematic way. In such simulations, stable configurations are hard to come by, and often require special fine-tuning. In addition, it is infeasible to run suites of well-resolved, realistic simulations with massive particles to drive planetary evolution where enough particles remain to represent the transneptunian populations to robustly statistically compare with observations. To complement state of the art full N-body simulations, we develop a method to artificially control each planet’s orbital elements independently from each other, which when carefully applied, can be used to test a wider suite of models. We modify two widely used publicly available N-body integrators: (1) the C code, REBOUND and (2) the FORTRAN code, Mercury. We show how the application of specific fictitious forces within numerical integrators can be used to tightly control planetary evolution to more easily explore migration and orbital excitation and damping. This tool allows us to replicate the impact a massive planetesimal disk would have on the planets, without actually including the massive planetesimals, thus decreasing the chaos and simulation runtime. We highlight an appropriate application that shows the impact of Neptune’s eccentricity damping and radial outward migration on Uranus’ eccentricity.
keywords:
keyword1 – keyword2 – keyword3
pubyear: 2024pagerange: Forcing Planets to Evolve: The Relationship Between Uranus and Neptune at Late Stages of Dynamical EvolutionA
@@ 1 Introduction
Planets’ orbits evolve through long and short term interactions with gas, planetesimals, and nearby planets (e.g., Chambers, 2009; Dawson & Johnson, 2018; Armitage, 2011; Andrews, 2020; Wyatt, 2008; Hughes et al., 2018; Pearce, 2024; Morbidelli & Levison, 2014; Levison & Dones, 2014; Morbidelli & Nesvorný, 2020; Horner et al., 2020; Fernandez & Ip, 1984; Goldreich et al., 2004). When modeling planetary evolution caused by gravitational interactions with millions of small bodies, simplifications are necessary to address an otherwise computationally infeasible problem. A common approach in simplifying and decreasing the computational cost of an N-body simulation is to artificially evolve a planet’s semimajor axis, eccentricity, and/or inclination. This is particularly important for science cases in which it is necessary to track the evolution of a statistically significant sample of planetesimals distinct from those driving the planetary evolution. For example, when studying the evolution of the outer solar system, artificially controlling the giant planets evolution is helpful so that the particles representing the Trans Neptunian Objects (TNOs) can be massless, increasing the number that can be modeled and tracked, and their final dynamical configuration can be compared to currently observed distributions and used to constrain evolution models.
@@
Artificial evolution has been widely done with "user-defined forces" where the code user defines an expression for the acceleration that will be added on top of the direct gravitational acceleration of the object due to other bodies in the simulation at each timestep (e.g., Malhotra, 1995; Hahn & Malhotra, 2005; Wolff et al., 2012; Lee & Peale, 2002; Ali-Dib et al., 2021). In celestial mechanics, force and acceleration are used interchangeably due to their proportional relationship for a given mass; for clarity, we will only refer to these as accelerations. In this paper we show how applying a "user-defined velocity" as well as a user-defined acceleration provides a complementary approach to what has been done before by granting stronger control over the bodies’ evolution, allowing each orbital element to be evolved independently from each other. Including user-defined velocities and accelerations in an N-body simulation mimics the presence of a physical process by updating the position and velocity of the particle at every time step. For example, one can emulate planets exchanging angular momentum with a disk of planetesimals by knowing how the orbital elements are expected to evolve with time for such a process. We extend the philosophy of user defined accelerations further by introducing user defined velocities into two widely used N-body integrators: (1) the publicly available C code, REBOUND (Rein & Liu, 2012; Rein & Tamayo, 2015; Rein et al., 2019) and (2) the FORTRAN code, Mercury6.2 (Chambers, 1999).
The use of artificial velocities and accelerations within numerical integrators is powerful when tight control of planets’ orbital evolution is desired. An example application is a chaotic evolution where minuscule changes to the initial parameters can produce a vastly different final configuration. Such chaos hinders the exploration of a large parameter space because the final outcomes are unpredictable. For example, in the well-studied planetary instability model of the outer solar system, a chaotic, planet-planet scattering phase is followed by a damping phase where the planetesimal disk is responsible for damping Neptune’s and Uranus’ eccentricities and migrating the planets outward. For this type of simulation, hundreds of runs are necessary just to get one that qualitatively matches the current planet configuration, making this a computationally expensive problem (e.g. Tsiganis et al., 2005; Levison et al., 2008; Clement et al., 2018; de Sousa et al., 2020). Individual control of the planets’ orbital elements is useful here in allowing exploration of a larger parameter space after the chaotic phase. While it is unphysical for the orbital elements to evolve independently (eccentricity and semimajor axis are coupled, for example), the use of this tool is appropriate when the orbits’ time evolution due to a physical process is known through analytical derivations and/or existing full N-body simulations. In this case, the time evolution of the eccentricity and semimajor axis of the planets due to angular momentum and energy exchange with the planetesimals is well-characterized by a function. Gravitational interactions between planetesimals and planets—Neptune in particular—sculpt the outer regions of the planetesimal disk over million and billion-year timescales. To rigorously compare simulations with observations, the final outcome of the planets must match their current observed orbits, where the possibility of such an evolution can be assessed by comparing the particle disk configuration to the dynamical structure of TNOs. Even in simulations where planet-planet scattering is not a major driver, planets often still need an extra push to arrive at the necessary location for detailed comparisons with observations.
To demonstrate an appropriate use of artificial orbital evolution, we focus on one class of frequently studied planet instability model: the simulation shown in Figure 1 of Tsiganis et al. (2005) (see Section 4). The planet-planet scattering phase produces unpredictable intermediate outcomes because of their chaotic nature, whereas Uranus’ and Neptune’s eccentricity damping phases are caused by dynamical friction. In addition, individual control of planet’s evolution is useful when one is interested in isolating the impact of gravitational interactions between specific bodies while avoiding other interactions that are also at play. An interesting outcome of evolving planets individually is noticing that eccentricity damping of one planet can cause another planet’s eccentricity to decline due to their secular coupling, as seen in Levison et al. (2011); Greenberg et al. (2013); Hansen & Murray (2015).
@@
In Section 2, we introduce the analytical framework for adding fictitious forces and velocities to the equations of motion and then describe modifications to the n-body codes Mercury 6.2 and REBOUND. In Section 3, we show examples with one and two planets where each planet’s orbital elements are evolved with arbitrary functions. In Section 4 we demonstrate two appropriate applications for when this code is useful. We show how we successfully reproduce the outward migration and eccentricity damping exhibited by Uranus and Neptune in a well-known Nice model simulation (Tsiganis et al., 2005). We also show that it is important to consider the secular coupling between Neptune and Uranus in the late-stages of orbital damping by planetesimals. As Neptune’s eccentricity damps, Uranus’ eccentricity damps indirectly as well. In Section 5 we summarise our results and discuss further applications for this code.
@@ 2 Modification to The Equations of Motion
We aim to adjust a simulated planet’s orbit such that, if no other perturbers were present in the simulation, each orbital element would independently evolve according to a specified function. The prescribed functions encapsulate the effects of processes not included in the simulation. Note that in simulations with multiple planets, since the planet is also accelerated by other system planets, the prescribed functions do not determine the full evolution of the orbit.
Additional forces are an existing functionality of REBOUND and Mercury6.2, where an arbitrary force can be added to a particle of choice every time-step, which operationally updates the particle’s velocity every time-step. However, it is not possible to fully and independently evolve a planet’s orbital elements using only additional forces. For full control, we require an update to the position of the planet at each time step (to account for changes to the velocity) as well as an update to the velocity. To calculate the required updates, we take partial derivatives of the equations of motion with respect to each orbital element to then form a fictitious velocity and acceleration to evolve the object given a functional form. This keeps the evolution of each orbital element independent from the others.
Physical accelerations such as radiation forces, gas dynamical friction, and more, are already incorporated onto REBOUND through the library REBOUNDx. The library contains eccentricity and inclination damping, but the only free parameter the user can choose is a damping timescale. With our modifications to REBOUND, rather than coding in individual physical effects, we can code arbitrary functional forms for how the elements should evolve. While this is nonphysical, these modifications allow us to probe a large amount of parameter space and various dynamical histories to learn how they impact the structure and dynamics of systems.
We follow Lee & Peale (2002) and expand on the work of Dawson & Murray-Clay (2012) and Wolff et al. (2012) to derive modifications to the equations of motion of an orbiting body to allow any form of evolution of the body’s orbital elements. Modifications to REBOUND and Mercury6.2 are outlined in Appendix A.
@@ 2.1 Analytical Derivation of Forces and Velocities for Arbitrary Orbital Element Evolution
$\dot{x}$ $=\dot{r}\cos{\Omega}\cos{(\omega+f)}-r\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (4a)
@@
$\dot{y}$ $=\dot{r}\sin{\Omega}\cos{(\omega+f)}+r\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (4b)
@@
$-r\sin{i}\dot{I}\cos{\Omega}\sin{(\omega+f)}-r\cos{i}\sin{\Omega}\dot{\Omega}\sin{(\omega+f)}$
@@
where we define $\dot{r}=\frac{\partial r}{\partial a}\dot{a}+\frac{\partial r}{\partial e}\dot{e}+\frac{\partial r}{\partial f}\dot{f}$. Recall that in the two body problem, $\dot{r}$ is derived assuming $a$ and $e$ are constant in time, producing $\dot{r}=\frac{\partial r}{\partial f}\dot{f}=\frac{nae\sin{f}}{\sqrt{1-e^{2}}}$ (see Murray & Dermott 1999).
@@
and similarly for $\dot{y}_{\rm user}$, $\ddot{y}_{\rm user}$, $\dot{z}_{\rm user}$, and $\ddot{z}_{\rm user}$ where $\dot{a}$, $\dot{e}$, $\dot{i}$, $\dot{\omega}$, and $\dot{\Omega}$ are the arbitrary functions we define (see Section 3). The user-defined velocity and user-defined acceleration are added to the existing motion of the particles. For each time step, $\Delta t$, the $x$-position and $x$-velocity of the object are updated by a term $\dot{x}_{\rm user}\Delta t$ and $\ddot{x}_{\rm user}\Delta t$, respectively and similarly for $y$ and $z$. All of the partial derivatives used in Equation (5) and further discussion of our implementation in REBOUND and Mercury6.2 can be found in Appendix A.
@@ 3 One and two planet examples
We demonstrate the potential of our modifications to REBOUND and Mercury6.2 with two example simulations, one with a star and one planet and the other with a star and two planets. For both examples, we ran 50 million year REBOUND (MERCURIUS integrator) and Mercury6.2 (HYBRID integrator) simulations. To compare short term and long term planetary behavior between the two N-body codes, we ran simulations with the same initial conditions, type of integrator, and output cadence. The five orbital elements, $a$, $e$, $i$, $\omega$, $\Omega$ are evolved using user-defined velocities and accelerations (Equation 5) by defining a functional form for $\dot{a}$, $\dot{e}$, $\dot{i}$, $\dot{\omega}$, and $\dot{\Omega}$. For the examples in this section, we choose a combination of arbitrary functional forms for the orbital elements shown in Equation (6), where $g$ is a stand-in variable for any of the evolved orbital elements:
@@
For the single-planet example, we integrate the orbits of a solar-mass star and Jupiter-mass planet. To confirm that each element is evolved independently from the others, we ran several initial tests (not shown) evolving one element at a time as well as combinations of elements. We find that the signature of one orbital element’s evolution is apparent on other elements to a negligible degree. For example, when exponentially or sinusoidally evolving only the eccentricity ($e$) of the planet, the semi-major axis ($a$) evolution contains a residual signature of the same functional form, though it is only noticeable at a precision of one part in $10^{7}$ for REBOUND and one part in $10^{4}$ for Mercury6.2. The perturbation on the semi-major axis is more pronounced when $\dot{e}$ in Equation (5) is large, confirming that the user-defined velocities and accelerations must remain small compared to the central gravitational acceleration.
@@
Figure 1 shows an example where we artificially evolve all orbital elements with arbitrary functions for $\dot{a}$, $\dot{e}$, $\dot{i}$, $\dot{\omega}$, and $\dot{\Omega}$. The Jupiter-mass planet is initialized at $a_{0}=5.2$ au, $e=0.2$, $i=10^{\circ}$, $\omega=50^{\circ}$, $\Omega=30^{\circ}$, and $f=240^{\circ}$. The parameters describing the artificial evolution of the planet are found in Table 1. Through the modifications described in Section 2 and Appendix A, we were able to successfully evolve all elements arbitrarily and independently from each other.
$g_{0}$ $\Delta g$ $\tau_{g}$ (Myr) Equation 7
$a$ 5.2 au 1.8 au 10 logarithmic (a)
$e$ 0.2 0.1 5 sinusoidal
$i$ 10 deg 5 deg 4 exponential (c)
$\omega$ 50 deg 35 deg 80 linear (d)
$\Omega$ 30 deg 60 deg 20 sinusoidal (b)
$f$ 240 deg
Table 1: Function and parameters for the one-planet example. Each orbital element’s time evolution is shown in Figure 1.
@@
Figure 1: Fifty million year evolution of the semimajor axis ($a$ (au); top left), eccentricity ($e$; top middle), inclination ($i$ (deg); top right), argument of pericenter ($\omega$ (deg); bottom left), longitude of ascending node ($\Omega$ (deg); bottom middle), and true anomaly ($f$ (deg); bottom right) of a Jupiter-mass planet orbiting a solar-mass star. The planet is initialized with the orbital elements in column $g_{0}$ in Table 1. The orbital elements $a$, $e$, $i$, $\omega$, and $\Omega$ are artificially time evolved by the arbitrary functions (dashed line) in Equation (7), denoted by the last column in Table 1. Each orbital element in fact evolves independently from each other and follows the functional form prescribed in the code.
@@
Figure 2: Fifty million year evolution of the semimajor axis ($a$), eccentricity ($e$), inclination ($i$) (orange) of a Jupiter-mass (top row) and Neptune-mass (bottom row) planet orbiting a solar-mass star. The orbital elements are initialized and artificially evolved by an exponential (dark maroon dashed line) with parameters described in Table 2. Similar to the one planet example in Figure 1, the additional user velocity and acceleration prescription evolves the orbital elements analogous to a desired function. However, the evolution of the eccentricity and inclination of each planet have a short timescale oscillation on top due to the secular effects between the planets. In contrast to the one planet example, where we can evolve the element to a final value, in this case the equation typically sets the minimum or maximum of the oscillation.
$g_{0}$ $\Delta g$ $\tau_{g}$ (Myr) Equation 7
$a$ 6 au $-1$ au 10 exponential (c)
$e$ 0.2 $-0.1$ 5 exponential (c)
planet 1 $i$ 5 deg $-3$ deg 20 exponential (c)
Jupiter-mass $\omega$ 50 deg
$\Omega$ 30 deg
$f$ 240 deg
$a$ 23 au 7.0 au 10 exponential (c)
$e$ 0.1 0.2 5 exponential (c)
planet 2 $i$ 10 deg $-8$ deg 20 exponential (c)
Neptune-mass $\omega$ 200 deg
$\Omega$ 280 deg
$f$ 250 deg
Table 2: Functions and parameters for the two-planet example orbiting a solar-mass star shown in Figure 2.
For the two-planet example shown in Figure 2, we integrate a Jupiter-mass planet and Neptune-mass planet orbiting a solar mass star for fifty million years. The planets are initialized at $a_{0}=6$ au, $e_{0}=0.2$, $i=5^{\circ}$, $\omega_{0}=50^{\circ}$, $\Omega_{0}=30^{\circ}$, $f_{0}=240^{\circ}$ and $a_{0}=23$ au, $e_{0}=0.1$, $i_{0}=10^{\circ}$, $\omega_{0}=200^{\circ}$, $\Omega_{0}=280^{\circ}$, $f_{0}=250^{\circ}$, respectively. We exponentially evolve their semimajor axis, eccentricity, and inclination with parameters described in Table 2. Including a second planet in the system introduces gravitational forces between the planets which were not present in the one-planet integration shown in Figure 1. In this three-body system, the planets secularly perturb each other, causing oscillations in the eccentricity and inclination with periods shorter than the damping timescale we prescribe. Rather than having full control of the final value of a given element, the scale of change, $\Delta g$ (i.e. final value minus initial value), in Equation (7c) sets an approximate final outcome in simulations with more than one planet. Jupiter’s secular effect on Neptune is greater than Neptune’s effect such that Neptune often over- or under-shoots the prescribed final eccentricity by an amount comparable to the amplitude of the secular oscillation. We find that the artificial exponential growth and decay of the eccentricity and inclination typically sets the final local maximum or minimum value.
The simulation setup in Figure 2 is such that Neptune’s secular inclination amplitude is large since Jupiter started at a high inclination. The secular eccentricity perturbations for both Jupiter and Neptune are small compared to the eccentricity user-defined perturbations (i.e. the total user input change is several times larger than the secular amplitude), whereas, the exponential change for the inclination is on the same order as the amplitude of the secular oscillations. Understanding how user-defined perturbations compare to the gravitational perturbations between planets is essential for running effective models and knowing the limit of control one has on the evolution of planets. We highlight this further in Figure 3 with three examples of Neptune’s inclination evolution with different values of $\Delta i$ and $\tau_{i}$ relative to the secular amplitude, $\mathcal{A}_{\rm sec}$ and secular oscillation timescale (1/frequency), $\tau_{\rm sec}$. When $\tau_{i}>\tau_{\rm sec}$, the long-term evolution is dominated by the user-defined forces. When $\Delta i>\mathcal{A}_{\rm sec}$, the prescribed sinusoidal evolution is noticeable. The three regimes are described in the figure caption.

### v2

# Forcing Planets to Evolve: Interactions Between Uranus and Neptune at Late Stages of Dynamical Evolution

@@ Abstract
In early Solar System numerical simulations, where chaos is a primary driver, it is difficult to explore parameter space in a systematic way. In such simulations, stable configurations are hard to come by, and often require special fine-tuning. In addition, it is infeasible to run suites of well-resolved, realistic simulations with a disk of massive particles to drive planetary evolution where enough particles remain to represent the transneptunian populations to robustly statistically compare with observations. To complement state of the art full N-body simulations, we develop a method to artificially control each planet’s orbital elements independently from each other, which when carefully applied, can be used to test a wider suite of models. We modify two widely used publicly available N-body integrators: (1) the C code, REBOUND and (2) the FORTRAN code, Mercury6.2. We show how the application of specific fictitious forces within numerical integrators can be used to tightly control planetary evolution to more easily explore migration and orbital excitation and damping. This tool allows us to replicate the impact a massive planetesimal disk would have on the planets, without actually including the massive planetesimals, thus decreasing the chaos and simulation runtime. We demonstrate the utility of this tool by applying it to the coupled orbital evolution of Uranus and Neptune, and show that Neptune’s eccentricity damping and radial outward migration have the appropriate affect on Uranus’ eccentricity.
planets and satellites: dynamical evolution and stability — methods: numerical
software: REBOUND: Rein & Liu (2012), NUMPY: van der Walt et al. (2011), MATPLOTLIB: Hunter (2007), PANDAS: Wes McKinney (2010), ASTROPY: Astropy Collaboration et al. (2013, 2018, 2022) , SCIPY: Virtanen et al. (2020)
@@ 1 Introduction
Planets’ orbits evolve through long and short term interactions with gas, planetesimals, and nearby planets (e.g., Fernandez & Ip, 1984; Lee & Peale, 2002; Goldreich et al., 2004; Chambers, 2009; Nesvorný, 2018; Morbidelli & Nesvorný, 2020). When modeling planetary evolution caused by gravitational interactions with millions of small bodies, simplifications are necessary to address an otherwise computationally infeasible problem. A common approach in simplifying and decreasing the computational cost of an N-body simulation is to artificially evolve a planet’s semimajor axis, eccentricity, and/or inclination. This is particularly important for science cases in which it is necessary to track the evolution of a statistically significant sample of planetesimals distinct from those driving the planetary evolution. For example, when studying the evolution of the outer solar system, artificially controlling the giant planets evolution is helpful so that the particles representing the Trans Neptunian Objects (TNOs) can be massless, increasing the number that can be modeled and tracked, and their final dynamical configuration can be compared to currently observed distributions and used to constrain evolution models (e.g Hahn & Malhotra, 2005; Levison et al., 2008; Dawson & Murray-Clay, 2012; Nesvorný, 2015; Nesvorný & Vokrouhlický, 2016).
@@
Artificial evolution has been widely done with “user-defined forces” where the code user defines an expression for the acceleration that will be added on top of the direct gravitational acceleration of the object due to other bodies in the simulation at each timestep (e.g., Malhotra, 1995; Hahn & Malhotra, 2005; Wolff et al., 2012; Lee & Peale, 2002; Ali-Dib et al., 2021). In celestial mechanics, the terms “force” and “acceleration” are used interchangeably due to their proportional relationship for a given mass; for clarity, we will only refer to these as accelerations. Additional accelerations are an existing functionality of widely used N-body integrators. The publicly available N-body codes REBOUND (Rein & Liu, 2012; Rein & Tamayo, 2015; Rein et al., 2019) and Mercury6.2 (Chambers, 1999), have the capacity to add an arbitrary acceleration to a particle of choice at every time-step, which operationally updates the particle’s velocity every time-step. This allows a user to model physical forces on the planet and make it move in a prescribed way. A number of useful physical accelerations such as radiation forces, gas dynamical friction, and more, are already incorporated onto REBOUND through the library REBOUNDx (Tamayo et al., 2020). The library also contains eccentricity and inclination damping, but the only free parameter the user can choose is a damping timescale. This combination of capabilities is enough to model physical forces on a particle. Neither code currently has the functionality to fully and independently evolve any combination of a planet’s orbital elements; such an evolution requires both a “user-defined velocity” and a “user-defined acceleration.” User-defined velocities are not typically included in these codes because it is non-physical for the orbital elements to change independently from each other; However, they have been employed by Lee & Peale (2002) to independently evolve semi-major axis and eccentricity and by Wolff et al. (2012) to independently evolve semi-major axis, eccentricity, and inclination.
In this paper we present modifications to REBOUND and Mercury6.2 that allow a user to evolve a particle’s orbital elements independently from each other because of user-defined accelerations and user-defined velocities. Our approach complements existing codes by allowing arbitrary functions to be chosen for how the elements should evolve. While such evolution may be nonphysical, these modifications allow us to probe a large amount of parameter space and number of dynamical histories motivated by more realistic models to learn how these evolutions impact the structure and dynamics of systems.
@@
An example context where criteria 1-3 are met is the well-known planetary instability model for the solar system (e.g. Tsiganis et al., 2005; Levison et al., 2008, 2011; Nesvorný, 2018; de Sousa et al., 2020), where the rearrangement and migration of the outer planets is typically ignited by Jupiter and Saturn crossing or exiting from a mutual mean motion resonance, which causes their orbital eccentricities to be excited. The resulting planetary instability excites the orbits of the other giant planets, and they undergo a chaotic era of strong planet-planet interactions for several millions to tens of millions of years. In simulations with successful outcomes that qualitatively match the real solar system, Uranus and Neptune, characterized by their eccentric and inclined orbits, are typically damped to more circular and coplanar orbits by dynamical friction (i.e. exchange of angular momentum and energy between the planets and nearby planetesimals). At this time, Jupiter and Saturn were most likely damped to some extent as well, but not as much as the outer ice giants. The details of when an instability occurred and how long or how far Neptune migrated remain uncertain.
This planetary instability model is characterized by a chaotic evolution where minuscule changes to the initial parameters can produce a vastly different final configuration. Such chaos hinders the exploration of a large parameter space because the final outcomes are unpredictable. For this type of simulation, hundreds of runs are necessary just to get one that qualitatively matches the current planet configuration, making this a computationally expensive problem (e.g. Tsiganis et al., 2005; Levison et al., 2008; Clement et al., 2018; de Sousa et al., 2020). Individual control of the planets’ orbital elements is useful here in allowing exploration of a larger parameter space after the chaotic phase. For example, instability models typically include some amount of planetesimal-driven migration and dynamical friction after the instability phase to push planets toward a configuration that is more similar to the current configuration of the solar system (see Section 4). Both of these physical processes are due to interactions between planetesimals and a planet, and their effect on the semimajor axis and eccentricity of that planet can be emulated by known functions in the simulation. Because the planets are no longer having direct, scattering interactions, we can exert more control to fictitiously evolve planets’ orbits and can explore parameter space. While it is unphysical for the orbital elements to evolve independently (eccentricity and semimajor axis are coupled, for example), fictitiously evolving planets’ orbits is appropriate when the orbits’ time evolution due to a physical process is known through analytical derivations and/or existing full N-body simulations. To rigorously compare these simulations with observations, the final outcome of the planets must match their current observed orbits, where the possibility of such an evolution can be assessed by comparing the particle disk configuration to the dynamical structure of TNOs. Even in simulations where planet-planet scattering is not a major driver, planets often still need an extra push to arrive at the necessary location for detailed comparisons with observations. Lastly, individual control of planet’s evolution is useful when one is interested in isolating the impact of gravitational interactions between specific bodies while avoiding other interactions that are also at play. An interesting outcome of evolving planets individually is noticing that eccentricity damping of one planet can cause another planet’s eccentricity to decline due to their secular coupling, as seen in Levison et al. (2011); Greenberg et al. (2013); Hansen & Murray (2015).
In Section 2, we introduce the analytical framework for adding fictitious forces and velocities to the equations of motion and then describe modifications to the n-body codes Mercury6.2 and REBOUND. In Section 3, we show examples with one and two planets where each planet’s orbital elements are evolved with arbitrary functions. In Section 4 we demonstrate two appropriate applications for when this code is useful. We show how we successfully reproduce the outward migration and eccentricity damping exhibited by Uranus and Neptune in a well-known Nice model simulation (Tsiganis et al., 2005). We also show that it is important to consider the secular coupling between Neptune and Uranus in the late-stages of orbital damping by planetesimals. Specifically, we demonstrate that as Neptune’s eccentricity damps, Uranus’ eccentricity damps indirectly as well. In Section 5 we summarise our results and discuss further applications for this code.
@@ 2 Modification to The Equations of Motion
We aim to adjust a simulated planet’s orbit such that, if no other perturbers were present in the simulation, each orbital element would independently evolve according to a specified function. The prescribed functions encapsulate the effects of processes not included in the simulation. Note that in simulations with multiple planets, since the planet is also accelerated by other system planets, the prescribed functions do not determine the full evolution of the orbit. It is not possible to fully and independently evolve a planet’s orbital elements using only additional forces. For full control, we require an update to the position of the planet at each time step (to account for changes to the velocity) as well as an update to the velocity. To calculate the required updates, we take partial derivatives of the equations of motion with respect to each orbital element to then form a fictitious velocity and acceleration to evolve the object given a functional form. This keeps the evolution of each orbital element independent from the others, which is not the case when only a user-defined acceleration is added. We discuss this detail further in Appendix B.1.
We follow Lee & Peale (2002) and expand on the work of Dawson & Murray-Clay (2012) and Wolff et al. (2012) to derive modifications to the equations of motion of an orbiting body to allow any form of evolution of the body’s orbital elements. Modifications to REBOUND and Mercury6.2 are outlined in Appendix B.3.
@@ 2.1 Analytical Derivation of Forces and Velocities for Arbitrary Orbital Element Evolution
$\dot{x}$ $=\dot{r}\cos{\Omega}\cos{(\omega+f)}-r\dot{\Omega}\sin{\Omega}\cos{(\omega+f)}$ (4a)
@@
$\dot{y}$ $=\dot{r}\sin{\Omega}\cos{(\omega+f)}+r\dot{\Omega}\cos{\Omega}\cos{(\omega+f)}$ (4b)
@@
$-r\dot{I}\sin{i}\cos{\Omega}\sin{(\omega+f)}-r\dot{\Omega}\cos{i}\sin{\Omega}\sin{(\omega+f)}$
@@
where we define $\dot{r}=\frac{\partial r}{\partial a}\dot{a}+\frac{\partial r}{\partial e}\dot{e}+\frac{\partial r}{\partial f}\dot{f}$.
Recall that in the two body problem, $\dot{r}$ is derived assuming $a$ and $e$ are constant in time, producing $\dot{r}=\frac{\partial r}{\partial f}\dot{f}=\frac{nae\sin{f}}{\sqrt{1-e^{2}}}$ (see Murray & Dermott 1999).
@@
and similarly for $\dot{y}_{\rm user}$, $\ddot{y}_{\rm user}$, $\dot{z}_{\rm user}$, and $\ddot{z}_{\rm user}$ where $\dot{a}$, $\dot{e}$, $\dot{I}$, $\dot{\omega}$, and $\dot{\Omega}$ are the arbitrary functions we define (see Section 3). The user-defined velocity and user-defined acceleration are added to the existing motion of the particles. For each time step, $\Delta t$, the $x$-position and $x$-velocity of the object are updated by a term $\dot{x}_{\rm user}\Delta t$ and $\ddot{x}_{\rm user}\Delta t$, respectively and similarly for $y$ and $z$. Further discussion of our implementation in REBOUND and Mercury6.2 can be found in Appendix B, including all of the partial derivatives used in Equation (5) (B.3).
@@ 3 One and two planet examples
We demonstrate the potential of our modifications to REBOUND and Mercury6.2 with example simulations consisting of one or two planets orbiting the Sun, using the MERCURIUS and HYBRID integrators for REBOUND and Mercury6.2, respectively. To compare short term and long term planetary behavior between the two N-body codes, we ran simulations with the same initial conditions, type of integrator, and output cadence. The five orbital elements, $a$, $e$, $i$, $\omega$, $\Omega$ are evolved using user-defined velocities and accelerations (Equation 5) by defining a functional form for $\dot{a}$, $\dot{e}$, $\dot{I}$, $\dot{\omega}$, and $\dot{\Omega}$. For the examples in this section, we choose a combination of arbitrary functional forms for the orbital elements shown in Equation (6), where $g$ is a stand-in variable for any of the evolved orbital elements:
@@
For the single-planet example, we integrate the orbit of a Jupiter-mass planet around a solar-mass star. To confirm that each element is evolved independently from the others, we ran several initial tests (not shown) evolving one element at a time as well as combinations of elements. We find that the signature of one orbital element’s evolution is apparent on other elements to a negligible degree. For example, when exponentially or sinusoidally evolving only the eccentricity ($e$) of the planet, the semi-major axis ($a$) evolution contains a residual signature of the same functional form, though it is only noticeable at a precision of one part in $10^{7}$ for REBOUND and one part in $10^{4}$ for Mercury6.2. The perturbation on the semi-major axis is more pronounced when $\dot{e}$ in Equation (5) is large enough for the additional force and velocity due to the eccentricity term are no longer small compared to the velocity and acceleration of the Keplerian motion. This confirms that the user-defined velocities and accelerations must remain small compared to the central gravitational acceleration.
@@
Figure 1 shows an example where we artificially evolve all orbital elements with arbitrary functions for $\dot{a}$, $\dot{e}$, $\dot{I}$, $\dot{\omega}$, and $\dot{\Omega}$. We simulate the Jupiter-mass planet’s motion for fifty million years. It is initialized with parameters and imposed evolution shown in Table 1. Through the modifications described in Section 2 and Appendix B, we were able to successfully evolve all elements arbitrarily and independently from each other.
@@
Figure 1: Fifty million year evolution of the semimajor axis ($a$ (au); top left), eccentricity ($e$; top middle), inclination (ii (deg); top right), argument of pericenter ($\omega$ (deg); bottom left), longitude of ascending node ($\Omega$ (deg); bottom middle), and true anomaly ($f$ (deg); bottom right) of a Jupiter-mass planet orbiting a solar-mass star. The planet is initialized with the orbital elements shown in Table 1. The orbital elements $a$, $e$, $i$, $\omega$, and $\Omega$ are artificially time evolved by the arbitrary functions (dashed line) in Equation (7), denoted by the “Eq” column in Table 1. Each orbital element in fact evolves independently from each other and follows the functional form prescribed in the code.
@@
Figure 2: One hundred fifty million year evolution of the semimajor axis ($a$), eccentricity ($e$), inclination (ii) (orange) of a Jupiter-mass (top row) and Neptune-mass (bottom row) planet orbiting a solar-mass star. The orbital elements are initialized and artificially evolved by an exponential (dark maroon dashed line) with parameters described in Table 1. Similar to the one planet example in Figure 1, the additional user velocity and acceleration prescription evolves the orbital elements analogous to a desired function. However, the evolution of the eccentricity and inclination of each planet have a short timescale oscillation on top due to the secular effects between the planets. In contrast to the one planet example, where we can evolve the element to a final value, in this case, the end result is near, but not exactly the final value. This is due to the two planets causing additional oscillations of each other’s orbital elements.